OpenStudy (anonymous):

can someone help with this question? X^2=101

OpenStudy (anonymous):

It is giving me the choices of square root 101 square root of 101, or negitive square root of 101 50.5 the solution is not a real number.

OpenStudy (anonymous):

\[x^2 = 101\] \[x=\sqrt{101}\] x=+/- 10.05

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

But I don't have that as a option

OpenStudy (anonymous):

what options do you have? could you post them?

OpenStudy (anonymous):

I posted them at the top of my post here

OpenStudy (anonymous):

squaroot of 101 is the answer

OpenStudy (anonymous):

yes i see =)

OpenStudy (anonymous):

pineda can you show me how you got that

OpenStudy (anonymous):

i already show you \[x^2=101\] \[x=\sqrt{101}\]

OpenStudy (anonymous):

x^2=101 squaroot bothsides to get rid of the squre of x lefting you answer x=√101 which can be read as squareroot of 101

OpenStudy (anonymous):

I just needed to see how it was done sorry and thank you